((e^x)+(4e^-x))/2 = 2
solution in terms of natural log and answer in decimal approx.?????

Answer by stanbon(75887)

((e^x)+(4e^-x))/2 = 2
solution in terms of natural log and answer in decimal approx.?????
----------------
e^x + [4/e^x] = 4
Multiply thru by e^x to get:
----------
e^(2x) + 4 = 4e^x
-----
Rearrange to quadratic form:
(e^x)^2 - 4e^x +4 = 0
----
Solve using quadratic formula.
e^x = [4 +- sqrt(16-4)]/2
---
e^x = [4+- 2sqrt(3)]/2
---
e^x = [2 +- sqrt(3)[
---
Separate:
e^x = 2+sqrt(3) or e^x = 2-sqrt(3
-----
x = ln(2+sqrt(3)) or x = ln(2-sqrt(3))
--------------
Approximation answers:
x = 1.3170 or x = -1.3170
================================
Cheers,
Stan H.
====================